Find the vector equation of the plane determined by the points (1, 0, −1),
(0, 1, 1) and (−1, 1, 0). Also nd the point of intersection of the line
1
Expert's answer
2020-02-10T09:34:34-0500

There are three points let's name them as

P(1,0,-1)

Q(0,1,1)

R(-1,1,0)

"\overrightarrow{PQ}=-i+j+2k \newline
\overrightarrow{RQ}=i+k"

cross product of these two vector will give us the normal of plane

"\begin{vmatrix}
 i & j &k \\
 -1 & 1&2\\1&0&1
\end{vmatrix}""=""i+""3j-k"

Equation of the vector can be written as

"a(x-x_0)+b(y-y_0)+c(z-z_0)=0"

"1(x-1)+3(y-0)-1(z-(-1))"

="x+3y-z-2=0"